Write the equation for the line parallel to y = 7x + 2 that passes through the point (7, -7).
Answer by feliz1965(151)

The equation we seek must have the same slope as the given equation. Lines with equal slopes are parallel.
The slope of y = 7x + 2 is the coefficient of x or 7.
We were given one point and now we know the slope.
Use the point-slope formula to find the equation.
y - y1 = m(x - x1)....This is the point-slope formula.
y - (-7) = 7(x - 7)
y + 7 = 7x - 49
y = 7x - 49 - 7
y = 7x - 56
Did you follow?